# Natural super-orbitals representation of many-body operators

### v1

@@ VI Conclusion
Acknowledgement. The author thank Serge Florens, Yuriel Núñez-Fernández and Frank Pollmann for interesting discussions. The author was supported by the Deutsche Forschungsgemeinschaft through the cluster of excellence ML4Q (EXC 2004, project-id 390534769) and by the Deutsche Forschungs- gemeinschaft through CRC 1639 NuMeriQS (project-id 511713970).
@@ Appendix A Weingarten calculation for Haar random unitaries
$\mathbb{E}_{U}(\mathcal{D})=\sum_{r=(\alpha,\beta)\in\mathrm{Rem}(\mathcal{D})}\mathcal{D}_{r}\mathrm{Wg}(d,\alpha\beta^{-1}),$ (38)
@@
Figure 9: Weingarten diagram for the trace of operators $\mathrm{Tr}\left(U^{\dagger}S^{z}_{i}US^{z}_{i}U^{\dagger}S^{z}_{i}US^{z}_{i}\right)$. We note $U^{\dagger}=\bar{U}^{t}$, where the transpose is performed by inverting the direct and dual spaces, denoted by empty and full circles respectively. Blue lines connect direct spaces of $U_{i}$ and $U_{\alpha(i)}$ and red lines connect the dual spaces of $U_{i}$ and $U_{\beta(i)}$. Solid and dashed lines correspond to the two different possible permutations for $\alpha$ and $\beta$.
@@
In the panel a) of Fig. 9, there is only a single choice for $\alpha$ and $\beta$, leading to a single reduced diagram. The reduced diagram reads $\mathcal{D}_{r}=\mathrm{Tr}(\hat{c}^{\dagger}_{m})\mathrm{Tr}(\hat{c}_{n})=0$, since these operator only have off-diagonal elements. The elements of $A$ are thus always 00, such that the eigenvalues of $R$ are trivially all equal to $1/2$.
@@ Appendix B Correlation spectrum in the atomic limit
$\begin{split}\left[\hat{U}(t)\right]&=\exp(itU\begin{bmatrix}0&0&0&0\\ 0&-\frac{1}{2}&\frac{V}{U}&0\\ 0&\frac{V}{U}&-\frac{1}{2}&0\\ 0&0&0&0\end{bmatrix})\\ &=e^{i\frac{\tilde{t}}{2}}\begin{bmatrix}e^{-i\frac{\tilde{t}}{2}}&0&0&0\\ 0&\cos(\omega)&i\sin(\omega)&0\\ 0&i\sin(\omega)&\cos(\omega)&0\\ 0&0&0&e^{-i\frac{\tilde{t}}{2}}\end{bmatrix},\end{split}$ (40)
where $\omega=V/U$ and $\tilde{t}=Ut$. We can compute the elements of the matrix $A$ defined as $[A]_{mn}=\mathrm{Tr}[\hat{U}^{\dagger}(t)\hat{c}^{\dagger}_{m}\hat{U}(t)\hat{c}_{n}]/\mathrm{Tr}[\mathds{1}]$, which reads:

### v2

@@ VI Conclusion and outlook
Acknowledgement. The author thank Serge Florens, Yuriel Núñez-Fernández, Xhek Turkeshi and Frank Pollmann for interesting discussions. The author was supported by the Deutsche Forschungsgemeinschaft through the cluster of excellence ML4Q (EXC 2004, project-id 390534769) and by the Deutsche Forschungs- gemeinschaft through CRC 1639 NuMeriQS (project-id 511713970).
@@ Appendix A Weingarten calculation for Haar random unitaries
$\mathbb{E}_{U}(\mathcal{D})=\sum_{r=(\alpha,\beta)\in\mathrm{Rem}(\mathcal{D})}\mathcal{D}_{r}\mathrm{Wg}(d,\alpha\beta^{-1}),\$(40)
@@
Figure 9: Weingarten diagram for the trace of operators $\mathrm{Tr}\left(\hat{O}\hat{c}^{\dagger}_{m}\hat{O}\hat{c}_{n}\right)$, corresponding to the elements of the matrix $A$ for Haar random unitaries. a) $\hat{O}=\hat{U}(t)$. b) $\hat{O}=\hat{U}^{\dagger}(t)(\hat{n}_{k}-1/2)\hat{U}(t)$. We note $U^{\dagger}=\bar{U}^{t}$, where the transpose is performed by inverting the direct and dual spaces, denoted by empty and full circles respectively. Blue lines connect direct spaces of $U_{i}$ and $U_{\alpha(i)}$ and red lines connect the dual spaces of $U_{i}$ and $U_{\beta(i)}$. Solid and dashed lines correspond to the two different possible permutations for $\alpha$ and $\beta$.
@@
In the panel a) of Fig. 9, there is only a single choice for $\alpha$ and $\beta$, leading to a single reduced diagram. The reduced diagram reads $\mathcal{D}_{r}=\mathrm{Tr}(\hat{c}^{\dagger}_{m})\mathrm{Tr}(\hat{c}_{n})=0$, since these operator only have off-diagonal elements. The elements of $A$ are thus always 0, such that the eigenvalues of $R$ are trivially all equal to $1/2$.
@@ Appendix B Correlation spectrum in the atomic limit
[…]
where $\omega=V/U$ and $\tilde{t}=Ut$. We can compute the elements of the matrix $A$ defined as $A]_{mn}=\mathrm{Tr}[\hat{U}^{\dagger}(t)\hat{c}^{\dagger}_{m}\hat{U}(t)\hat{c}_{n}]/\mathrm{Tr}[\mathds{1}]$, which reads: